OpenStudy (anonymous):

write the expression as a complex number in standard form identify the real and imaginary parts (-11+3i)+(4-6i)

OpenStudy (anonymous):

When adding complex number of the form: \[a+b\iota\] you need to add up the real and imaginary components separately. In this case we have: \[(-11+3\iota)+(4-6\iota)=-7-3\iota\] i,e,: \[-11+4=-7\] and \[3-6=-3\] Then we can identify the real and imaginary parts.